OpenStudy (anonymous):

[6.02] and [6.03] Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

what method do u want to use?

OpenStudy (anonymous):

x-y=10 +2x+y=2 --------- 3x = 12 which solve and get x =4, then you plug 4 into one of the equations and solve for y and get y = -6